# I Introduction

### v2

@@ Acknowledgements.
The Fermi-LAT Collaboration acknowledges support for LAT development, operation and data analysis from NASA and DOE (United States), CEA/Irfu and IN2P3/CNRS (France), ASI and INFN (Italy), MEXT, KEK, and JAXA (Japan), and the K.A. Wallenberg Foundation, the Swedish Research Council and the National Space Board (Sweden). Science analysis support in the operations phase from INAF (Italy) and CNES (France) is also gratefully acknowledged. This work performed in part under DOE Contract DE-AC02-76SF00515. P.V. acknowledges support from NASA grant NNM11AA01A. M.M. acknowledges support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program Grant agreement No. 948689 (AxionDM). L.B. acknowledges support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy—EXC 2121 “Quantum Universe”. P.D.V. acknowledge “funding by the European Union – NextGenerationEU” RFF M4C2 project IR0000012 CTA+. We acknowledge financial support from INAF through the “Ricerca Fondamentale 2024” program (mini-grant titled ‘Measurement of the intergalactic magnetic field with gamma-ray sources’ (PI A.S.).
Supplemental Material
CRPropa Simulations Details
@@
Figure 5: Comparison of pair-echo SEDs for $\log_{10}(B/\mathrm{G})=-20$ and a log-parabola VHE intrinsic spectrum, using different EBL models (see text for details)
@@
We assume an IGMF with a turbulent zero-mean Gaussian random field and a Kolmogorov spectrum. It is defined in Fourier space, transformed into real space, and then projected onto a uniformly spaced cubic grid with $N=100^{3}$ cells of 50 Mpc size. The minimum and maximum scale lengths in which the magnetic field are defined are 1 Mpc and 25 Mpc, respectively, resulting in a correlation length of $\ell_{B}\sim 5\,$Mpc. The cubic cell is then periodically repeated to cover the distance between the source and Earth. We conduct the simulations for different values of the root mean square of the IGMF field strength between $10^{-20}\,$G and $10^{-15}\,$G. For each tested field and VHE spectrum with different cutoff energies, we inject $2\times 10^{5}$ photons. For the two strongest IGMF strengths ($\log_{10}(B/\mathrm{G})=-15.5$ and 15-15), we inject $4\times 10^{5}$ and $8\times 10^{5}$ photons, respectively, due to the strong time dilution of the pair-echo signal. All particles are traced with a step size of at least 10-5 pc which enable us to reproduce time delays with an accuracy better than 20 minutes.
To evaluate the pair-echo SED within a certain time interval $\Delta T$, we consider an observer perfectly aligned with the jet axis and select all cascade photons arriving with a time delay $\Delta T$ [25]. In the left panel of Fig. 4, we show an example of the SED derived in the time interval 3.3-365 days for the three different VHE intrinsic spectra tested; a log parabola and a log parabola multiplied with two different exponential cutoff energies.
In principle we also need to consider the spatial extension of the cascade. However, we have verified that for all the tested IGMF strenghts and time delays (up to 1 year) the cascade signal is well within the point spread function (PSF) of the Fermi LAT (see right panel of Fig. 4), so that we can consider the cascade emission as point like.
@@
Fermi-LAT Analysis Details
@@
The background is fitted by (1) optimizing the RoI by means of a standard iterative process [^1] https://fermipy.readthedocs.io/en/latest/fitting.html#id2, (2) removing those components with a preliminary detection TS <4<4 or a predicted number of counts lower than 3, and (3) re-fitting the normalization of all sources within $3^{\circ}$ of GRB 221009A. For the time bin from $T_{0}+3.3$ d to one year, the final likelihood (for each IGMF strength) is obtained by including an additional point-like source whose spectrum is described with results from the CRPropa cascade simulations. For the other time bins we consider three distinct scenarios:
@@
    Cascade only: A single source with the simulated cascade spectrum is added to the RoI. The spectrum is fixed to that obtained in the simulations (both normalisation and shape).
@@
    Phenomenological afterglow with cascade: In addition to the simulated cascade, an additional component for a point-like source is added to account for the GeV emission from the GRB. Its spectrum is described with a power-law with $\Gamma=2$ [39] and, in contrast to the cascade component, its normalization is left free in the fit.
@@
    Physical afterglow with cascade: In addition to the simulated cascade, an additional component for a point-like source is added to account for the GeV emission, whose normalization and spectral shape is also fixed to that predicted by the afterglow model during the corresponding time interval.
The likelihood profiles for different scenarios are presented in Fig. 6, with the complete list of limits derived for each scenario summarized in Tab. 1. Additionally, we present the SEDs for a point-like source at the position of GRB 221009A with and without a power-law component extracting the GeV signal (Fig. 8). Fluxes for the afterglow component are fully consistent with those derived in [39].
Figure 6: Likelihood profiles for further cases. Top panels display the likelihood profiles obtained by not extracting any signal (left) and by subtracting an afterglow component according the a physical model (right). Both cases assume an exponential cut-off at 7 TeV. Bottom panels display the profiles derived assuming an ad-hoc afterglow emission with $\Gamma=2$, but with the cut-off at 13 TeV (left) or without it (right). Solid lines represent cubic spline interpolations between the derived likelihood values.
$T_{0}$-$T_{f}$ Cut-off at 7 TeV Cut-off at 13 TeV No cut-off
Power-law GRB model Power-law GRB model Power-law GRB model
@@
Table 1: The lower bounds on the IGMF strength, expressed in G at the 95% confidence level, are presented for all the tested time intervals. These results assume the power-law and GRB physically-motivated models as spectral models for the GRB in the Fermi-LAT analysis. The results were obtained using the log-parabola model, Eq. (1). Exponential cut-offs are considered by including an additional term to the equation.
To estimate if similar studies would further improve the constraints by increasing the exposure time, we extrapolate the derived exposure map of the ROI up to 5 years. A $2\sigma$ confidence level sensitivity curve is derived, requiring at least 3 counts from the echo-emission and considering the impact of diffuse components. We derive such sensitivity curves for narrow energy bands, with four bins per energy decade and assuming a flat spectrum with $\Gamma=2$ within each energy bin. We note that these narrow bands are required to account for the energy-dependent time evolution of the cascade. Additional CRPropa simulations up to $T_{0}+5$ yr predict a measurable (yet moderate) improvement in the sensitivity from 3 to 15 GeV (Fig. 9), the most favorable energy range for the Fermi-LAT.
Figure 7: Left: SED cascades for the different time bins considered, in comparison with the LAT observations during the corresponding time window. Right: Same SED cascades in comparison with the upper limits derived from LAT observations once the afterglow emission was subtracted with a power-law component during the corresponding time window.
@@
SSC afterglow model at GeV energies
@@
We numerically solve for the bulk Lorentz factor of the blast-wave, $\Gamma$, as a function of radius and follow the distribution of the electrons’ random Lorentz factor, $\gamma_{e}$ through the continuity equation. We assume the electrons develop a power-law distribution as they are injected into the blast wave from the ISM. We take into account radiative and adiabatic cooling.
To obtain the spectral energy distribution, at each time we convolve the synchrotron and SSC emissivity with the electron distribution. To compare the model with the LHAASO measurements, we integrate the flux over their reported time range. We vary the free parameters of the model (total energy $E_{0}$, intial Lorentz factor $\Gamma_{0}$, electron distribution index $s$, electron $\xi_{e}$ and magnetic field $\xi_{B}$ equipartition parameter) to match the observed LHAASO lightcurve (Tab. 2). We do not perform a fit. Once we achieve an acceptable match to the LHAASO light curve, we integrate the model over the required time bin and derive the expected SED for the LAT energy range (Fig. 10).
@@
Figure 10: SEDs for the SSC model during the time bins defined in Table 3 of [39]. Each color represents a different time bin. The two components correspond to synchrotron and SSC emission (dotted and dashed-dotted lines, respectively). Dashed lines represent the total emission once EBL absorption is considered. The LAT energy range, in green, is used in our likelihood analysis for the physical afterglow with cascade scenario.

### v3

@@ Acknowledgements.
The Fermi-LAT Collaboration acknowledges support for LAT development, operation and data analysis from NASA and DOE (United States), CEA/Irfu and IN2P3/CNRS (France), ASI and INFN (Italy), MEXT, KEK, and JAXA (Japan), and the K.A. Wallenberg Foundation, the Swedish Research Council and the National Space Board (Sweden). Science analysis support in the operations phase from INAF (Italy) and CNES (France) is also gratefully acknowledged. This work performed in part under DOE Contract DE-AC02-76SF00515. P.V. acknowledges support from NASA grant NNM11AA01A. M.M. acknowledges support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program Grant agreement No. 948689 (AxionDM). L.B. acknowledges support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy—EXC 2121 “Quantum Universe”. P.D.V. acknowledges “funding by the European Union – NextGenerationEU” RFF M4C2 project IR0000012 CTA+. We acknowledge financial support from INAF through the “Ricerca Fondamentale 2024” program (mini-grant titled ‘Measurement of the intergalactic magnetic field with gamma-ray sources’ (PI A.S.).
Appendix A CRPropa Simulations Details
@@
Figure 5: Comparison of pair-echo SEDs for $\log_{10}(B/\mathrm{G})=-20$ and a log-parabola VHE intrinsic spectrum, using different EBL models (see text for details).
@@
We assume an IGMF with a turbulent zero-mean Gaussian random field and a Kolmogorov spectrum. It is defined in Fourier space, transformed into real space, and then projected onto a uniformly spaced cubic grid with $N=100^{3}$ cells of 50 Mpc size. The minimum and maximum scale lengths in which the magnetic field are defined are 1 Mpc and 25 Mpc, respectively, resulting in a correlation length of $\ell_{B}\sim 5\,$Mpc. The cubic cell is then periodically repeated to cover the distance between the source and Earth. We conduct the simulations for different values of the root mean square of the IGMF field strength between $10^{-20}\,$G and $10^{-15}\,$G. For each tested field and VHE spectrum with different cutoff energies, we inject $2\times 10^{5}$ photons. For the two strongest IGMF strengths ($\log_{10}(B/\mathrm{G})=-15.5$ and 15-15), we inject $4\times 10^{5}$ and $8\times 10^{5}$ photons, respectively, due to the strong time dilution of the pair-echo signal. All particles are traced with a step size of at least 10-5 pc which enables us to reproduce time delays with an accuracy better than 20 minutes.
To evaluate the pair-echo SED within a certain time interval $\Delta T$, we consider an observer perfectly aligned with the jet axis [^2] The assumption was justified by existing estimates, see Table 3 in [57]. and select all cascade photons arriving with a time delay $\Delta T$ [17]. In the left panel of Fig. 4, we show an example of the SED derived in the time interval 3.3-365 days for the three different VHE intrinsic spectra tested; a log parabola and a log parabola multiplied with two different exponential cutoff energies.
In principle we also need to consider the spatial extension of the cascade. However, we have verified that, for all of the tested IGMF strengths and time delays (up to one year), the cascade signal is well within the point spread function (PSF) of the Fermi LAT (see right panel of Fig. 4), so that we can consider the cascade emission as point like.
@@
Appendix B Fermi-LAT Analysis Details
@@
The background is fitted by (1) optimizing the RoI by means of a standard iterative process [^3] https://fermipy.readthedocs.io/en/latest/fitting.html#id2, (2) removing those components with a preliminary detection TS <4<4 or a predicted number of counts lower than 3, and (3) re-fitting the normalization of all sources within $3^{\circ}$ of GRB 221009A. For the time bin from $T_{0}+3.3$ d to one year, the final likelihood (for each IGMF strength) is obtained by including an additional pointlike source whose spectrum is described with results from the CRPropa cascade simulations. For the other time bins we consider three distinct scenarios:
@@
    Cascade only: A single source with the simulated cascade spectrum is added to the RoI. The spectrum is fixed to that obtained in the simulations (both normalization and shape).
@@
    Phenomenological afterglow with cascade: In addition to the simulated cascade, an additional component for a pointlike source is added to account for the GeV emission from the GRB. Its spectrum is described with a power law with $\Gamma=2$ [9] and, in contrast to the cascade component, its normalization is left free in the fit.
@@
    Physical afterglow with cascade: In addition to the simulated cascade, an additional component for a pointlike source is added to account for the GeV emission, whose normalization and spectral shape is also fixed to that predicted by the afterglow model during the corresponding time interval.
The likelihood profiles for different scenarios are presented in Fig. 6, with the complete list of limits derived for each scenario summarized in Tab. 1. Additionally, we present the SEDs for a pointlike source at the position of GRB 221009A with and without a power law component extracting the GeV signal (Fig. 8). Fluxes for the afterglow component are fully consistent with those derived in [9].
Figure 6: Likelihood profiles for further cases. Top panels display the likelihood profiles obtained by not extracting any signal (left) and by subtracting an afterglow component according to the physical model in Appendix C (right). Both cases assume an exponential cutoff at 7 TeV. Bottom panels display the profiles derived assuming an ad hoc afterglow emission with $\Gamma=2$, but with the cutoff at 13 TeV (left) or without it (right). Solid lines represent cubic spline interpolations between the derived likelihood values.
$T_{0}$-$T_{f}$ cutoff at 7 TeV cutoff at 13 TeV No cutoff
power law GRB model power law GRB model power law GRB model
@@
Table 1: The lower bounds on the IGMF strength, expressed in G at the 95% confidence level, are presented for all the tested time intervals. These results assume the power law and GRB physically motivated models as spectral models for the GRB in the Fermi-LAT analysis. The results were obtained using the log-parabola model, Eq. (1). Exponential cutoffs are considered by including an additional term to the equation.
To estimate if similar studies would further improve the constraints by increasing the exposure time, we extrapolate the derived exposure map of the ROI up to 5 years. A $2\sigma$ confidence level sensitivity curve is derived, requiring at least 3 counts from the echo emission and considering the impact of diffuse components. We derive such sensitivity curves for narrow energy bands, with four bins per energy decade and assuming a flat spectrum with $\Gamma=2$ within each energy bin. We note that these narrow bands are required to account for the energy-dependent time evolution of the cascade. Additional CRPropa simulations up to $T_{0}+5$ yr predict a measurable (yet moderate) improvement in the sensitivity from 3 to 15 GeV (Fig. 9), the most favorable energy range for the Fermi-LAT.
Figure 7: Left: SED cascades for the different time bins considered, in comparison with the LAT observations during the corresponding time window. Right: Same SED cascades in comparison with the upper limits derived from LAT observations once the afterglow emission was subtracted with a power law component during the corresponding time window.
@@
Appendix C SSC afterglow model at GeV energies
@@
We numerically solve for the bulk Lorentz factor of the blast wave, $\Gamma$, as a function of radius and follow the distribution of the electrons’ random Lorentz factor, $\gamma_{e}$ through the continuity equation. We assume the electrons develop a power law distribution as they are injected into the blast wave from the interstellar medium. We take into account radiative and adiabatic cooling.
To obtain the spectral energy distribution, at each time we convolve the synchrotron and SSC emissivity with the electron distribution. To compare the model with the LHAASO measurements, we integrate the flux over their reported time range. We vary the free parameters of the model (total energy $E_{0}$, initial Lorentz factor $\Gamma_{0}$, electron distribution index $s$, electron $\xi_{e}$ and magnetic field $\xi_{B}$ equipartition parameter) to match the observed LHAASO light curve (Tab. 2). We do not perform a fit. Once we achieve an acceptable match to the LHAASO light curve, we integrate the model over the required time bin and derive the expected SED for the LAT energy range (Fig. 10).
@@
Figure 10: SEDs for the SSC model during the time bins defined in Table 3 of [9]. Each color represents a different time bin. The two components correspond to synchrotron and SSC emission (dotted and dash-dotted lines, respectively). Dashed lines represent the total emission once EBL absorption is considered. The LAT energy range, in green, is used in our likelihood analysis for the physical afterglow with cascade scenario.